# Towards Robust and Realible Multimodal Fake News Detection with Incomplete Modality

### v1

@@ Abstract.
Multimodal fake news detection (MFND) has become an urgent task with the emergence of huge multimodal fake content on social media platforms. Previous studies mainly focus on complex feature extraction and fusion to learn discriminative information from multimodal content. However, in real-world applications, multimedia news may naturally lose some information during dissemination, resulting in modality incompleteness, which is detrimental to the generalization and robustness of existing models. To this end, we propose a novel generic and robust multimodal fusion strategy, termed Multi-expert Modality-incomplete Learning Network (MMLNet), which is simple yet effective. It consists of three key steps: (1) Multi-Expert Collaborative Reasoning to compensate for missing modalities by dynamically leveraging complementary information through multiple experts. (2) Incomplete Modality Adapters compensates for the missing information by leveraging the new feature distribution. (3) Modality Missing Learning leveraging an label-aware adaptive weighting strategy to learn a robust representation with contrastive learning. We evaluate MMLNet on three real-world benchmarks across two languages, demonstrating superior performance compared to state-of-the-art methods while maintaining relative simplicity. By ensuring the accuracy of fake news detection in incomplete modality scenarios caused by information propagation, MMLNet effectively curbs the spread of malicious misinformation. Code is publicly available at https://github.com/zhyhome/MMLNet.
Incomplete Modality, Robust Fake News Detection, Social Media
Figure 1. Illustration of fake news scenarios with incomplete modalities. Existing works have only investigated the scenario of complete modalities in the first step of the distortion propagation theory.:
1. Introduction
@@
With the widespread use of social media platforms like Twitter and Weibo, users can easily share personal opinions, emotions, and daily information. However, the misuse of these platforms and the lack of effective regulation have led to the rapid spread of misinformation, making the issue of fake news increasingly severe (Zubiaga et al., 2018). This phenomenon has profound social impacts and poses significant threats to areas such as politics, finance, and public health. The abuse of advanced generative models has further exacerbated the problem, contributing to the rise of text-based fake news (Sheng et al., 2021) and visual deepfakes (Zhou et al., 2021). In addition, the dissemination of multimodal forged media has increased the reach of false information, increasing its potential to mislead the public (Liu et al., 2024). Given the coexistence of fake information in both the image and text modalities, detecting such fake news presents an unprecedented challenge.
@@
To address this challenge, current multimodal fake news detection (MFND) methods (Liu et al., 2024; Tong et al., 2024; Liu et al., 2025; Wu et al., 2025), primarily focus on the fusion of cross-modal features. Despite the significant achievements of these methods, they do not consider scenarios involving incomplete modalities due to missing modality information in real-world settings, which introduces risks for practical applications.
@@
According to the Communication Distortion Theory, during the process of information transmission, the information may undergo distortion, misunderstanding, or loss, resulting in incomplete or inaccurate information. Early studies (Melumad et al., 2021) have demonstrated this theory through behavioral experiments. (Zhang et al., 2024b, 2025b) also demonstrates that early detection in the propagation path helps identify misinformation.
@@
Figure 1 visually illustrates the application of this theory in the context of news dissemination. In Stage 1, multi-media initially publishes the original news. However, after passing through Group Users 1, the information may become distorted. For simplicity, we consider only the case of information loss. As a result, Group Users 2 will receive incomplete news, and different user groups will obtain news with varying degrees of completeness. Unfortunately, existing MFND methods are limited to verifying the authenticity of news at the first stage, relying on complete news data in both training and testing while overlooking the fact that multiple user groups receive incomplete multimodal news information. This limitation poses risks in real-world applications. For the first time, we address this issue and propose the Multi-expert Modality-incomplete Learning Network (MMLNet), which mitigates this phenomenon by incorporating incomplete multimodal data during both training and testing.
Although existing MFND approaches have achieved significant success, their performance degrades substantially in highly incomplete modality scenarios. This decline occurs because these methods overly emphasize multimodal fusion representations while neglecting the joint representation of unimodal information. Consequently, when unimodal information is missing, it further disrupts the effectiveness of multimodal fusion representations. We conduct comparative experiments on complete and incomplete modalities using the latest publicly released state-of-the-art open-sourced MFND methods and finetuned advanced Multimodal Large Language Models (MLLM), including NSLM (Dong et al., 2024), MIMoE (Liu et al., 2025), Qwen-VL (Bai et al., 2023), and LLaVA (Liu et al., 2023), on real-world datasets in two different languages, as shown in Table 1.
In multimodal learning, extensive research has been conducted on the handling of incomplete modalities (Han et al., 2022; Sun et al., 2022; Yuan et al., 2021; Wu et al., 2024; Xu et al., 2024). Existing approaches in incomplete multimodal learning (Yuan et al., 2021; Wu et al., 2024; Xu et al., 2024) primarily focus on reconstructing missing modalities from available ones (Li et al., 2024b; Zhang et al., 2024a), aiming to maintain feature consistency across samples with arbitrary modality gaps. While these techniques have shown promising results, their direct application to image-text pair based MFND presents two fundamental challenges. First, existing methods are typically developed independently for different problems, leading to heterogeneous models with distinct design strategies. This architectural divergence complicates direct integration and increases the alignment burden. Second, misinformation in social media often results in semantic inconsistencies between images and text, which contradicts the widely held assumption in domains such as video and medical imaging (Lian et al., 2023; Pham et al., 2019; Wang et al., 2023) that modalities originating from the same sample are inherently correlated. This lack of cross-modal coherence hinders the effectiveness of conventional approaches in MFND, necessitating novel solutions tailored to its unique challenges.
The most related work is (Zeng et al., 2025), however, which focuses on detecting the authenticity of video modalities and is coarse-grained for the absence of incomplete modalities (missing the entire modality), which is not suitable for the scenario studied in this paper.
To fill this gap, we first investigate the impact of incomplete modalities on MFND and propose MMLNet. The main contributions of this paper can be summarized as follows:
  • We focus on robust multimodal fake news detection under incomplete modalities, an aspect that has been less emphasized in previous research but is prevalent in real-world scenarios. To the best of our knowledge, this is a pioneering exploration of model robustness in MFND.
  • We propose MMLNet, a novel approach to handle arbitrary incomplete modality information for MFND tasks in a unified framework. It leverages multi-expert learning of feature distributions of various modalities, compensates for missing modality information through incomplete modality adapters, and further employs modality missing learning to enhance modality distribution for robust MFND.
  • Experimental results show that MMLNet achieves significant improvements over sota MFND methods and MLLM on three real-world MFND datasets across two languages, validating its superiority in robust MFND.
@@
Table 1. Part of experimental results show that the Accuracy score (%) of existing SOTA models has declined in the face of incomplete modalities. Incomplete modalities are dominated by incomplete text modalities and incomplete image modalities in this case. Note that ‘-’ indicates that the modality is complete, and ‘↓’ indicates the percentage of modality missing. Green fonts show the performance degradation of the model compared to the complete modality scenario.
@@
↓75(%) ↓25(%) MMLNet(ours) 92.55↓2.66 80.19↓7.59
2. Related Work
2.1. Fake News Detection
2.1.1. Unimodal Methods
@@
Text-based approaches focus on different analytical dimensions. Some methods integrate social context by leveraging propaganda strategies (Huang et al., 2023), user feedback (Min et al., 2022), temporal trends (Hu et al., 2023), and news environments (Sheng et al., 2022). Ghanem et al. (Ghanem et al., 2021) proposed a technique that combines topic and sentiment features extracted from textual data. More recently, Nan et al. (Nan et al., 2021) and Zhu et al. (Zhu et al., 2023) explored domain shift issues caused by variations in word frequency and sentiment, introducing domain gating and domain memory mechanisms to improve domain representation.
@@
Various image-based methods (Yang et al., 2022) have been introduced to assess the authenticity of visual content by analyzing edited traces. Some CNN-based techniques operate in the spatial domain to identify artifact patterns, utilizing approaches such as blending (Li et al., 2020a), multiple instance learning (Li et al., 2020b), patch consistency analysis (Zhao et al., 2021), reconstruction techniques (Huang et al., 2020; Liang et al., 2022), and local feature mining (Dong et al., 2022). Alternatively, other studies focus on transforming images into the frequency domain through methods like discrete cosine transform (Qian et al., 2020), composite phase spectrum analysis (Liu et al., 2021), and high-frequency noise extraction (Luo et al., 2021).
2.1.2. Multimodal Methods
@@
Mainstream multimodal MFND methods extract semantic representations by integrating cross-modal textual and visual features (Jaiswal et al., 2017). Some approaches enhance semantic information by incorporating external knowledge bases, as proposed by Sabir et al. (Sabir et al., 2018) and Wang et al. (Wang et al., 2020). Qi et al. (Qi et al., 2021) focused on extracting visual entities to improve high-level news comprehension. Peng et al. (Jin et al., 2017) applied named entity recognition and word segmentation to enhance Chinese social media identification. To better fuse textual and visual features, co-attention networks (Wu et al., 2021) and contextual attention networks (Qian et al., 2021) have been developed. Kochkina et al. (Zubiaga et al., 2017) employed a multi-task learning approach for rumor verification. Additionally, Ying et al. (Ying et al., 2023) introduced an improved Mixture of Experts network to refine feature integration across multiple perspectives. To mitigate spurious correlations from modality inconsistencies and data biases, ambiguity learning (Chen et al., 2022) and causal inference (Chen et al., 2023) were introduced separately. FAK-Owl (Liu et al., 2024) performs knowledge-enhanced verification based on Large Vision-Language Models. NSLM (Dong et al., 2024) utilizes neuro-symbolic reasoning for misinformation detection. MIAN (Zhang et al., 2025a) is an attention network based on the utilization of intrinsic discriminative features. DGA-FAKE (Zhang et al., 2024b) detects fake news early through the diffusion propagation path. Recent work MIMoE (Liu et al., 2025) detects misinformation through interactive mixture of experts.
@@
These methods effectively utilize image and text modalities to uncover implicit misinformation but struggle to handle missing modalities. Since such issues frequently occur in real-world applications (Melumad et al., 2021), we propose MMLNet, a robust and practical approach for detecting misinformation in news.
2.2. Robust Multimodal Learning
@@
The issue of missing modalities is prevalent across all types of multimodal data and has been widely studied. Early research employed data imputation methods (Lian et al., 2023) to recover missing modalities from available ones. Some approaches directly fill in missing modalities with fixed values (Chen and Zhang, 2020; Zhang et al., 2022). Additionally, some researchers leverage the generative capabilities of specific neural network architectures, such as autoencoders (Vincent et al., 2008) and Transformers (Vaswani et al., 2017). (Hu et al., 2024; Guo et al., 2024; Lee et al., 2023) mitigate incomplete modalities through prompt learning. Some recovery-based methods (Li et al., 2024a; Xu et al., 2024; Wu et al., 2024; Zhang et al., 2024a; Huan et al., 2024), have been proven to be effective in robust multimodal learning.
@@
However, these modality recovery-based methods are not well-suited for MFND, as they inherently rely on modality consistency. Misinformation is often embedded in modality inconsistencies, making modality recovery detrimental to the model’s performance. Therefore, we utilize the modal feature distribution learned from multiple perspectives by MMLNet, make up for the missing features with the invisible feature distribution through the incomplete modality adapter, and effectively enhance the robustness of the modal features through modality missing learning.
3. Method
3.1. Problem Definition
3.1.1. Multimodal Fake News Detection
@@
The MFND task is a typical binary classification problem, where the positive class corresponds to fake news and the negative class corresponds to real news. Given a multi-modal news dataset $\mathcal{D}$, where $\mathcal{D}=\{x_{i},y_{i}\mid x_{i}=(x_{i}^{T},x_{i}^{V})\}_{i=1}^{N}.$ $x_{i}^{T}$represents textual information, $x_{i}^{V}$represents visual information, and $y_{i}$ denotes the corresponding label. To effectively distinguish fake news, the model’s objective is to learn the probability distribution of $\mathbb{P}(y\mid x)$.
3.1.2. Incomplete Modality
@@
Figure 2. Overview of MMLNet framework.:
3.2. Overview Framework
@@
3.3. Multi-Expert Reasoning Module
3.3.1. Expert of Text-Modality
In multimodal learning, the text modality typically plays a dominant role due to its higher information density. An incomplete text modality significantly impacts model performance; therefore, modeling text modality experts is crucial for the MFND task.

### v2

@@ Abstract
Multimodal Misinformation Recognition has become an urgent task with the emergence of huge multimodal fake content on social media platforms. Previous studies mainly focus on complex feature extraction and fusion to learn discriminative information from multimodal content. However, in real-world applications, multimedia news may naturally lose some information during dissemination, resulting in modality incompleteness, which is detrimental to the generalization and robustness of existing models. To this end, we propose a novel generic and robust multimodal fusion strategy, termed Multi-expert Modality-incomplete Learning Network (MMLNet), which is simple yet effective. It consists of three key steps: (1) Multi-Expert Collaborative Reasoning to compensate for missing modalities by dynamically leveraging complementary information through multiple experts. (2) Incomplete Modality Adapters compensates for the missing information by leveraging the new feature distribution. (3) Modality Missing Learning leveraging an label-aware adaptive weighting strategy to learn a robust representation with contrastive learning. We evaluate MMLNet on three real-world benchmarks across two languages, demonstrating superior performance compared to state-of-the-art methods while maintaining relative simplicity. By ensuring the accuracy of misinformation recognition in incomplete modality scenarios caused by information propagation, MMLNet effectively curbs the spread of malicious misinformation. Code is publicly available at https://github.com/zhyhome/MMLNet.
1 Introduction
@@
With the widespread use of social media platforms like Twitter and Weibo, users can easily share personal opinions, emotions, and daily information. However, the misuse of these platforms and the lack of effective regulation have led to the rapid spread of misinformation, making the issue of misinformation increasingly severe survey_rumours . This phenomenon has profound social impacts and poses significant threats to areas such as politics, finance, and public health. The abuse of advanced generative models has further exacerbated the problem, contributing to the rise of text-based fake news text_fake and visual deepfakes visual_fake . In addition, the dissemination of multimodal forged media has increased the reach of false information, increasing its potential to mislead the public fkaowl . Given the coexistence of fake information in both the image and text modalities, detecting such misinformation presents an unprecedented challenge.
@@
To address this challenge, current Multimodal Misinformation Recognition (MMR) methods fkaowl , MOE_WWW25 , MMDFND , 2025-efficient-COLING primarily focus on the fusion of cross-modal features. Despite the significant achievements of these methods, they do not consider scenarios involving incomplete modalities due to missing modality information in real-world settings, which introduces risks for practical applications.
Figure 1: Illustration of misinformation scenarios with incomplete modalities. Existing works have only investigated the scenario of complete modalities in the first step of the distortion propagation theory.
@@
According to the Communication Distortion Theory, during the process of information transmission, the information may undergo distortion, misunderstanding, or loss, resulting in incomplete or inaccurate information. Early studies data_missing_theory have demonstrated this theory through behavioral experiments. DGAmm24 , 10750410 also demonstrates that early recognition in the propagation path helps identify misinformation.
@@
Figure 1 visually illustrates the application of this theory in the context of news dissemination. In Stage 1, multi-media initially publishes the original news. However, after passing through Group Users 1, the information may become distorted. For simplicity, we consider only the case of information loss. As a result, Group Users 2 will receive incomplete news, and different user groups will obtain news with varying degrees of completeness. Unfortunately, existing MMR methods are limited to verifying the authenticity of news at the first stage, relying on complete news data in both training and testing while overlooking the fact that multiple user groups receive incomplete multimodal news information. This limitation poses risks in real-world applications. For the first time, we address this issue and propose the Multi-expert Modality-incomplete Learning Network (MMLNet), which mitigates this phenomenon by incorporating incomplete multimodal data during both training and testing.
@@
Table 1: Part of experimental results show that the Accuracy score (%) of existing SOTA models has declined in the face of incomplete modalities. Incomplete modalities are dominated by incomplete text modalities and incomplete image modalities in this case. Note that ‘-’ indicates that the modality is complete, and ‘↓’ indicates the percentage of modality missing. Green fonts show the performance degradation of the model compared to the complete modality scenario.
@@
[…]
Although existing MMR approaches have achieved significant success, their performance degrades substantially in highly incomplete modality scenarios. This decline occurs because these methods overly emphasize multimodal fusion representations while neglecting the joint representation of unimodal information. Consequently, when unimodal information is missing, it further disrupts the effectiveness of multimodal fusion representations. We conduct comparative experiments on complete and incomplete modalities using the latest publicly released state-of-the-art open-sourced MMR methods and finetuned advanced Multimodal Large Language Models (MLLM), including NSLM NSLMaaai24 , MIMoE MOE_WWW25 , Qwen-VL Qwen-VL , and LLaVA liu2023llava , on real-world datasets in two different languages, as shown in Table 1.
In multimodal learning, extensive research has been conducted on the handling of incomplete modalities MMAlign_emnlp22 , Sun2022EfficientMT , MM24robustSenDistribution , MM24robustSenMOE , robustSenMM21 . Existing approaches in incomplete multimodal learning MM24robustSenDistribution , MM24robustSenMOE , robustSenMM21 primarily focus on reconstructing missing modalities from available ones nips24robustSen1 , recover4 , aiming to maintain feature consistency across samples with arbitrary modality gaps. While these techniques have shown promising results, their direct application to image-text pair based MMR presents two fundamental challenges. First, existing methods are typically developed independently for different problems, leading to heterogeneous models with distinct design strategies. This architectural divergence complicates direct integration and increases the alignment burden. Second, misinformation in social media often results in semantic inconsistencies between images and text, which contradicts the widely held assumption in domains such as video and medical imaging GCNetTPAMI , translationAAAI19 , nips23robustEmo that modalities originating from the same sample are inherently correlated. This lack of cross-modal coherence hinders the effectiveness of conventional approaches in MMR, necessitating novel solutions tailored to its unique challenges.
The most related work is IMOL , however, which focuses on detecting the authenticity of video modalities and is coarse-grained for the absence of incomplete modalities (missing the entire modality), which is not suitable for the scenario studied in this paper.
To fill this gap, we first investigate the impact of incomplete modalities on MMR and propose MMLNet. The main contributions of this paper can be summarized as follows:
  • We focus on robust multimodal misinformation recognition under incomplete modalities, an aspect that has been less emphasized in previous research but is prevalent in real-world scenarios. To the best of our knowledge, this is a pioneering exploration of model robustness in MMR.
  • We propose MMLNet, a novel approach to handle arbitrary incomplete modality information for MMR tasks in a unified framework. It leverages multi-expert learning of feature distributions of various modalities, compensates for missing modality information through incomplete modality adapters, and further employs modality missing learning to enhance modality distribution for robust MMR.
  • Experimental results show that MMLNet achieves significant improvements over sota MMR methods and MLLM on three real-world MMR datasets across two languages, validating its superiority in robust MMR.
2 Related Work
2.1 Misinformation recognition
2.1.1 Unimodal Methods
@@
Text-based approaches focus on different analytical dimensions. Some methods integrate social context by leveraging propaganda strategies huang-etal-2023-faking , user feedback 10.1145/3485447.3512163 , temporal trends hu-etal-2023-learn , and news environments sheng-etal-2022-zoom . Ghanem et al. ghanem-etal-2021-fakeflow proposed a technique that combines topic and sentiment features extracted from textual data. More recently, Nan et al. weibo21 and Zhu et al. 10.1109/TKDE.2022.3185151 explored domain shift issues caused by variations in word frequency and sentiment, introducing domain gating and domain memory mechanisms to improve domain representation.
@@
Various image-based methods Yang2022ConfidenceCalibratedFI have been introduced to assess the authenticity of visual content by analyzing edited traces. Some CNN-based techniques operate in the spatial domain to identify artifact patterns, utilizing approaches such as blending Li_2020_CVPR_fake , multiple instance learning 10.1145/3394171.3414034 , patch consistency analysis Zhao_2021_ICCV_fake , reconstruction techniques 10.1145/3394171.3413732 , 10.1007/978-3-031-19781-9_8 , and local feature mining Dong2022ImplicitIL . Alternatively, other studies focus on transforming images into the frequency domain through methods like discrete cosine transform Qian2020ThinkingIF , composite phase spectrum analysis Liu2021SpatialPhaseSL , and high-frequency noise extraction Luo2021GeneralizingFF .
2.1.2 Multimodal Methods
@@
Mainstream multimodal MMR methods extract semantic representations by integrating cross-modal textual and visual features 10.1145/3123266.3123385 . Some approaches enhance semantic information by incorporating external knowledge bases, as proposed by Sabir et al. 10.1145/3240508.3240707 and Wang et al. 10.1145/3372278.3390713 . Qi et al. 10.1145/3474085.3481548 focused on extracting visual entities to improve high-level news comprehension. Peng et al. weibo applied named entity recognition and word segmentation to enhance Chinese social media identification. To better fuse textual and visual features, co-attention networks wu-etal-2021-multimodal and contextual attention networks 10.1145/3404835.3462871 have been developed. Kochkina et al. pheme employed a multi-task learning approach for rumor verification. Additionally, Ying et al. 10.1609/aaai.v37i4.25670 introduced an improved Mixture of Experts network to refine feature integration across multiple perspectives. To mitigate spurious correlations from modality inconsistencies and data biases, ambiguity learning 10.1145/3485447.3511968 and causal inference chen-etal-2023-causal were introduced separately. FAK-Owl fkaowl performs knowledge-enhanced verification based on Large Vision-Language Models. NSLM NSLMaaai24 utilizes neuro-symbolic reasoning for misinformation recognition. yin2024fine proposed a method based on heterogeneous graphs for graph attention networks, achieving fine-grained multimodal deepfake classification. MIAN zhang2025multimodalinverseattentionnetwork is an attention network based on the utilization of intrinsic discriminative features. DGA-FAKE DGAmm24 detects misinformation early through the diffusion propagation path. Recent work MIMoE MOE_WWW25 detects misinformation through interactive mixture of experts.
@@
These methods effectively utilize image and text modalities to uncover implicit misinformation but struggle to handle missing modalities. Since such issues frequently occur in real-world applications data_missing_theory , we propose MMLNet, a robust and practical approach for detecting misinformation in news.
2.2 Robust Multimodal Learning
@@
The issue of missing modalities is prevalent across all types of multimodal data and has been widely studied. Early research employed data imputation methods GCNetTPAMI to recover missing modalities from available ones. Some approaches directly fill in missing modalities with fixed values 10.1145/3394486.3403182 , 9258396 . Additionally, some researchers leverage the generative capabilities of specific neural network architectures, such as autoencoders 10.1145/1390156.1390294 and Transformers 10.5555/3295222.3295349 . guo-etal-2024-multimodal , hu2024deep , lee2023cvpr mitigate incomplete modalities through prompt learning. Some recovery-based methods recover5 , recover1 , MM24robustSenDistribution , MM24robustSenMOE , recover4 have been proven to be effective in robust multimodal learning. Furthermore, researchers in the relevant field have proposed methods to enhance the robustness of multimodal learning. cai2025multi proposed a causal-guided adaptive multimodal diffusion network to address the displacement phenomenon that may occur when samples from different source domains interfere with each other during the learning process. NORM-TR liu2025noise reduces the impact of noise on multimodal learning through a noise perception learning method.
Figure 2: Overview of MMLNet framework.
@@
However, these modality recovery-based methods are not well-suited for MMR, as they inherently rely on modality consistency. Misinformation is often embedded in modality inconsistencies, making modality recovery detrimental to the model’s performance. Therefore, we utilize the modal feature distribution learned from multiple perspectives by MMLNet, make up for the missing features with the invisible feature distribution through the incomplete modality adapter, and effectively enhance the robustness of the modal features through modality missing learning.
3 Method
3.1 Problem Definition
3.1.1 Multimodal Misinformation recognition
@@
The MMR task is a typical binary classification problem, where the positive class corresponds to fake news and the negative class corresponds to real news. Given a multi-modal news dataset $\mathcal{D}$, where $\mathcal{D}=\{x_{i},y_{i}\mid x_{i}=(x_{i}^{T},x_{i}^{V})\}_{i=1}^{N}.$ $x_{i}^{T}$represents textual information, $x_{i}^{V}$represents visual information, and $y_{i}$ denotes the corresponding label. To effectively distinguish fake news, the model’s objective is to learn the probability distribution of $\mathbb{P}(y\mid x)$.
3.1.2 Incomplete Modality
@@
3.2 Overview Framework
@@
3.3 Multi-Expert Reasoning Module
3.3.1 Expert of Text-Modality
In multimodal learning, the text modality typically plays a dominant role due to its higher information density. An incomplete text modality significantly impacts model performance; therefore, modeling text modality experts is crucial for the MMR task.